OpenStudy (anonymous):

Solve the linear equation (X+6)/2-1=(x+6)/7

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Compassionate

OpenStudy (phi):

\[ \frac{(x+6)}{2}-1=\frac{(x+6)}{7} \] ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

multiply both sides by 14 (which is the common denominator of 2 and 7) when you multiply the left side be sure to multiply 14 times *both* terms can you do that ?

OpenStudy (anonymous):

No I can't, I don't get it

OpenStudy (phi):

to multiply (without simplifying) write 14 * next to each term can you do that ?

OpenStudy (anonymous):

Yes, one moment please

OpenStudy (anonymous):

14(x+6)/2-1=(x-6)14/7

OpenStudy (phi):

ok but you forgot the 1

OpenStudy (anonymous):

I have to multiply 4 by one too?

OpenStudy (phi):

yes, if you have 1+1 =2 and you multiply both sides by 14 you do 14*1 + 14*1 = 2*14 which (if you work it out) is still true. if you just did 14*1 + 1 = 2*14 it's gone horribly wrong.

OpenStudy (anonymous):

Is it -8?

OpenStudy (phi):

so you should have 14(x+6)/2-14*1=(x-6)14/7

OpenStudy (phi):

\[ \frac{14(x+6)}{2} - 14 = \frac{14(x-6)}{7} \\ \frac{14}{2}(x+6) - 14 = \frac{14}{7}(x-6) \] now simplify 14/2 on the left side and 14/7 on the right side can you do that?

OpenStudy (anonymous):

I already figured out the answer thanks